### v1

# Intrinsic nonlinear Hall effect beyond Bloch geometry

Abstract
The theory of the intrinsic Hall effect, both linear and nonlinear, is rooted in a geometry which is defined in the Bloch-vector parameter space; the formal expressions are mostly derived from semiclassical concepts. When disorder and interaction are considered there is no Bloch vector to speak of; one needs a more general quantum geometry, defined in a different parameter space. Such higher-level geometrical formulation of the intrinsic Hall effect provides very compact expressions, which have the additional virtue—in the Bloch special case—of yielding the known results in a straightforward way: the logic is not concealed by the algebra.
Introduction.–The second-order dc conductivity tensor is comprised of three terms which—in a semiclassical treatment—scale as the zeroth, first, and second power of the transport lifetime $\tau$. The first two terms are of the Hall kind, i.e. the induced current is normal to the electric field, and their main entries are quantum-geometrical entities: Berry connections and curvatures in Bloch space [1, 2, 3, 4, 5]. The main focus of the present work is on the $\tau^{0}$ term, which—to the best of the author’s knowledge—was first discovered in Ref. [1], where it was attributed to the field-induced “positional shift” of Bloch electrons; it will be indicated as $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$ in the following. This term has received much attention in recent times [6, 7, 8, 9, 10, 11].
I adopt here an exact quantum-mechanical framework, where by “exact” I mean that I address a system of interacting electrons and static classical nuclei, possibly in a disordered configuration. No relaxation time $\tau$ can enter the theory: in fact the concept itself does not make sense beyond the semiclassical theory. In the present framework the response functions are causal but nondissipative: therefore the induced current grows with some powers of time. At first order the intrinsic Hall current is $t^{0}$, and the Drude current is $t^{1}$ (the electrons undergo free acceleration). At second order one expects three terms, namely $t^{0}$, $t^{1}$, and $t^{2}$, in one-to-one correspondence to the three semiclassical terms; while for two of them a many-body formulation exists [12, 13], the $t^{0}$ term remained elusive so far. Its expression is provided here; for the sake of completeness the other terms are also addressed.
The parameter space of conventional quantum geometry is defined by the Bloch vector $\bf k$}, and the state vectors are the cell-periodic $\bf k$}-dependent Bloch orbitals; the observables are Fermi-volume integrals. Here instead I adopt the many-body formulation of quantum geometry [14, 15, 16, 17, 18], where the role of the Bloch vector is played instead by the “flux” $\kappa$ entering the kinetic term of the many-body Hamiltonian. The main entry of $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$ is cast here as a positional-shift tensor: a Berry curvature of an hybrid kind, whose variables are the flux and the electric field. Geometrical tensors of this kind may appear exotic in the context of conductivities; instead they are the main entries in the theory of polarization (formerly called “modern”) [19] and of other observables [20].
Many-body quantum geometry addresses in principle even systems with disorder and correlation, but has the additional virtue that it allows for compact and very transparent notations; all geometrical formulaæ can be easily converted when needed—e.g. for DFT implementation—in their more prolix Bloch counterparts. This is shown here in detail for the three terms of second order conductivity.
Exact theory.–The flux-dependent many body Hamiltonian has been introduced in 1964 by Kohn [21]; when modified in order to account for the absence of time-reversal (T) symmetry it reads
$\hat{H}=\frac{1}{2m}\sum_{i=1}^{N}\left[{\bf p}_{i}+\frac{e}{c}{\bf A}({\bf r}_{i})+\hbar{\mbox{$\kappa$}}\right]^{2}+\hat{V}$. (1)
It addresses a system of NN $d$-dimensional electrons in a cubic box of volume $L^{d}$. The flux $\kappa$ is a vector potential cast in inverse-length units, constant in space; the potential $\hat{V}$ includes the one-body potential (possibly disordered) and electron-electron interaction.
The system is macroscopically homogeneous; the $\kappa$-dependent eigenstates $|\Psi_{n}\rangle$ are normalized to one in the hypercube of volume $L^{Nd}$. The thermodynamic limit $N\rightarrow\infty$, $L\rightarrow\infty$, $N/L^{d}$ constant is understood throughout this work; the $\kappa$-derivatives must be evaluated first, and the $L\rightarrow\infty$ limit taken afterwards [21, 22, 23].
Periodic boundary conditions (PBCs) are assumed: the many-body wavefunctions are periodic with period $L$ over each electron coordinate ${\bf r}_{i}$ independently; the potential $\hat{V}$ enjoys the same periodicity. The vector potential in Eq. (1) summarizes all intrinsic T-breaking terms, as e.g. those due to a coupling to a background of local moments; even ${\bf A}({\bf r})$ obeys PBCs. The vector potential could even account for a macroscopic $\bf B$} field, provided that it is commensurate, and that the PBCs are modified accordingly [24]. A time-independent $\kappa$ amounts to a pure gauge transformation, but—as shown by Kohn—owing to PBCs the gauge-invariance is broken: the eigenvalues $E_{n}$ depend on $\kappa$ in the metallic case, while they are $\kappa$-independent in insulators [21, 22, 23, 25].
Kohn’s Hamiltonian has the virtue that the macroscopic current density can be cast as
$\hat{\bf j}=-\frac{e}{L^{d}}\hat{{\bf v}},\qquad\hat{{\bf v}}=\frac{1}{\hbar}\partial_{{\mbox{$\kappa$}}}\hat{H}$; (2)
furthermore a flux linear in time corresponds to perturbing the system with the dc field
[…]
$\frac{\partial j_{\alpha}(t)}{\partial{\cal E}_{\beta}}=\frac{e^{2}}{\hbar L^{d}}\left[-\Omega(\kappa_{\alpha},\kappa_{\beta})+\frac{t}{\hbar}\frac{\partial^{2}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}}\right],$ (4)
where $\Omega(\kappa_{\alpha},\kappa_{\beta})=-2\,\mbox{Im }\langle\partial_{k_{\alpha}}\Psi_{0}|\partial_{k_{\beta}}\Psi_{0}\rangle$ is the many-body Berry curvature [16].
The linear conductivity obtains by setting ${\mbox{$\kappa$}}=0$; switching from the $t$-domain to the $\omega$-domain and enforcing causality the second term in Eq. (4) yields the familiar linear Drude conductivity
$\sigma^{(\rm Drude)}_{\alpha\beta}(\omega)=\frac{D_{\alpha\beta}}{\pi}\frac{i}{\omega+i\eta}=D_{\alpha\beta}\left[\delta(\omega)+\frac{i}{\pi\omega}\right],$ (5)
where $\eta=0^{+}$ and
$D_{\alpha\beta}=\frac{\pi e^{2}}{\hbar^{2}L^{d}}\frac{\partial^{2}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}}$ (6)
evaluated at ${\mbox{$\kappa$}}=0$ is known as the Drude weight [21, 22, 23].
Both $\Omega(\kappa_{\alpha},\kappa_{\beta})$ and $D_{\alpha\beta}$ acquire a time dependence when expanded to the next order in $\kappa$:
$\partial_{t}\Omega(\kappa_{\alpha},\kappa_{\beta})$ $=$ $\partial_{\kappa_{\gamma}}\Omega(\kappa_{\alpha},\kappa_{\beta})\dot{\kappa}_{\gamma},$
$\partial_{t}D_{\alpha\beta}$ $=$ $\frac{\pi e^{2}}{\hbar^{2}L^{d}}\frac{\partial^{3}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}\partial\kappa_{\gamma}}\dot{\kappa}_{\gamma}.$ (7)
Therefore, owing to Eqs. (3) and (4), one has
$\frac{\partial^{2}j_{\alpha}(t)}{\partial{\cal E}_{\beta}\partial{\cal E}_{\gamma}}$ $=$ $\frac{\partial^{2}j_{\alpha}(0)}{\partial{\cal E}_{\beta}\partial{\cal E}_{\gamma}}+\frac{e^{3}t}{\hbar^{2}L^{d}}\partial_{k_{\gamma}}\Omega(\kappa_{\alpha},\kappa_{\beta})$ (8)
- $\frac{\pi e^{3}t^{2}}{\hbar^{3}L^{d}}\frac{\partial^{3}E_{0}}{\partial\kappa_{\alpha}\partial\kappa_{\beta}\partial\kappa_{\gamma}},$
where all derivatives are evaluated at ${\mbox{$\kappa$}}=0$
The three terms are parsed by their $t$-dependence: $t^{0}$, $t^{1}$, anf $t^{2}$, respectively. The $t^{2}$ term accounts for the quadratic Drude conductivity, and coincides with the expression found by Watanabe and Oshikawa [12]; the $t^{1}$ term is the many-body formulation of the nonlinear Hall effect previously established by the present author [13]. The $t^{0}$ term is the main focus of this work: as said above, it will be indicated as $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$.
[…]
$\langle\Psi_{0}|\,\hat{r}_{\alpha}\,|\Psi_{n}\rangle=i\langle\Psi_{0}|\partial_{k_{\alpha}}\Psi_{n}\rangle=-i\hbar\frac{\langle\Psi_{0}|\,\hat{v}_{\alpha}\,|\Psi_{n}\rangle}{{E_{0}-E_{n}}}.$ (10)
[…]
${\bf P}^{(\rm el)}=-\frac{ie}{L^{d}}\langle\Psi_{0}|\partial_{{\mbox{$\kappa$}}}\Psi_{0}\rangle,\quad{\mbox{$\kappa$}}=0,$ (11)
[…]
${\cal G}_{\alpha\gamma}$ $=$ $\frac{1}{e}\partial_{{\cal E}\gamma}{\cal A}_{\alpha}=\frac{i}{e}\langle\partial_{{\cal E}\gamma}\Psi_{0}|\partial_{k_{\alpha}}\Psi_{0}\rangle+\mbox{c.c.}$ (12)
$=$ $\frac{i}{e}(\,\langle\partial_{{\cal E}\gamma}\Psi_{0}|\partial_{k_{\alpha}}\Psi_{0}\rangle-\langle\partial_{k_{\alpha}}\Psi_{0}|\partial_{{\cal E}\gamma}\Psi_{0}\rangle\,)$
$=$ $\frac{1}{e}\Omega({\cal E}_{\gamma},\kappa_{\alpha}).$
${\cal G}_{\alpha\gamma}$ is therefore cast as an hybrid geometrical tensor of the same family as those entering polarization theory [19]. In fact it is shown elsewhere [20] that the curvature $\Omega({\cal E}_{\gamma},\kappa_{\alpha})$, evaluated at ${\mbox{$\kappa$}}=0$, yields the linear polarizability of an insulator:
$\chi_{\alpha\gamma}=-\frac{e}{L^{d}}\Omega({\cal E}_{\gamma},\kappa_{\alpha}).$ (13)
Notice that $\Omega({\cal E}_{\gamma},\kappa_{\alpha})$ is antisymmetric for the exchange ${\cal E}_{\gamma}\leftrightarrow\kappa_{\alpha}$, and is symmetric for the exchange $\gamma\leftrightarrow\alpha$.
[…]
$\mbox{Im }\sigma_{\alpha\gamma}(\omega)=\frac{D_{\alpha\gamma}}{\pi\omega}+\mbox{Im }\sigma^{(\rm regular)}_{\alpha\gamma}(\omega),$ (14)
where the regular term is linear in $\omega$; in insulators the expression is the same but with a vanishing Drude term. Is it then easy to show that in both insulators and metals one has
$\lim_{\omega\rightarrow 0}\frac{\mbox{Im }\sigma^{(\rm regular)}_{\alpha\gamma}(\omega)}{\omega}=\frac{e}{L^{d}}\Omega(\kappa_{\alpha},{\cal E}_{\gamma}).$ (15)
In metals this regular term is not observable because it is obliterated by the divergent imaginary part of $\sigma_{\alpha\beta}^{(\rm Drude)}(\omega)$; it manifests itself at second order via its $\kappa$-derivatives.
By evaluating ${\cal G}_{\alpha\gamma}$ in the parallel-transport gauge one gets (again both in metals and insulators)
${\cal G}_{\alpha\gamma}$ $=$ $-2\,\mbox{Re }\sum_{n\neq 0}\frac{\langle\Psi_{0}|\,\hat{r}_{\alpha}\,|\Psi_{n}\rangle\langle\Psi_{n}|\,\hat{r}_{\gamma}\,|\Psi_{0}\rangle}{E_{0}-E_{n}}$ (16)
$=$ $-2\hbar^{2}\,\mbox{Re }\sum_{n\neq 0}\frac{\langle\Psi_{0}|\,\hat{v}_{\alpha}\,|\Psi_{n}\rangle\langle\Psi_{n}|\,\hat{v}_{\gamma}\,|\Psi_{0}\rangle}{(E_{0}-E_{n})^{3}}.$
Finally, the sought for expression for the positional-shift conductivity is
$\sigma^{\rm(ps)}_{\alpha\beta\gamma}=\frac{e^{3}}{\hbar L^{d}}(\,\partial_{k_{\alpha}}{\cal G}_{\beta\gamma}-\partial_{k_{\beta}}{\cal G}_{\alpha\gamma}\,);$ (17)
even this expression is evaluated at ${\mbox{$\kappa$}}=0$. Eqs. (16) and (17) are reminiscent of the corresponding Bloch expressions in the literature: Eqs. (22) and (23) below.
Kohn-Sham theory.–In the special case of a crystalline system of noninteracting electrons the $L\rightarrow\infty$ limit can be performed analytically thanks to translational symmetry: all intensive observables are then expressed as Fermi-volume integrals.
In a Kohn-Sham framework the ground-state $|\Psi_{0}\rangle$ is a Slater determinant of single-particle orbitals. Whenever an operator $\hat{O}$ is the sum of one-body operators $\tilde{O}$ its expectation value is given by the sum of the expectation values of $\tilde{O}$ over the occupied orbitals. Here the orbitals are the Bloch orbitals $|\psi_{j{\bf k}}\rangle={\rm e}^{i{\bf k}\cdot{\bf r}}|u_{j{\bf k}}\rangle$ with eigenvalues $\epsilon_{j{\bf k}}$, normalized in the crystal cell of volume $V_{\rm cell}$—they are normalized differently from $|\Psi_{0}\rangle$. The formulæ below are given per spin channel; for spinless electrons in jargon.
Quantum geometry deals with the $|u_{j{\bf k}}\rangle$, which are eigenstates of the Kohn-Sham $\kappa$-dependent Hamiltonian
$H_{\bf k}=\frac{1}{2m}\left({\bf p}+\frac{e}{c}{\bf A}({\bf r})+\hbar{\bf k}+\hbar{\mbox{$\kappa$}}\right)^{2}+V({\bf r}),$ (18)
hence a $\kappa$-derivative evaluated at ${\mbox{$\kappa$}}=0$ coincides with a $\bf k$}-derivative. The one-body version of the curvature of Eq. (12) is then
$G_{j\alpha\gamma}=\frac{i}{e}(\,\langle\partial_{{\cal E}\gamma}u_{j{\bf k}}|\partial_{k_{\alpha}}u_{j{\bf k}}\rangle-\langle\partial_{k_{\alpha}}u_{j{\bf k}}|\partial_{{\cal E}\gamma}u_{j{\bf k}}\rangle\,),$ (19)
and the positional-shift conductivity is
$\sigma^{\rm(ps)}_{\alpha\beta\gamma}=\frac{e^{3}}{\hbar V_{\rm cell}}\sum_{j{\bf k}}\theta(\epsilon_{\rm F}-\epsilon_{j{\bf k}})(\,\partial_{k_{\alpha}}G_{j\beta\gamma}-\partial_{k_{\beta}}G_{j\alpha\gamma}\,),$ (20)
where $\epsilon_{\rm F}$ is the Fermi energy. In the $L\rightarrow\infty$ limit
$\frac{1}{V_{\rm cell}}\sum_{{\bf k}}\rightarrow\int_{\rm BZ}\frac{d{\bf k}}{(2\pi)^{d}}\;,$ (21)
$\sigma^{\rm(ps)}_{\alpha\beta\gamma}=\frac{e^{3}}{\hbar}\sum_{j}\int_{\rm BZ}\frac{d{\bf k}}{(2\pi)^{d}}\;\theta(\epsilon_{\rm F}-\epsilon_{j{\bf k}})(\,\partial_{k_{\alpha}}G_{j\beta\gamma}-\partial_{k_{\beta}}G_{j\alpha\gamma}\,).$ (22)
This is identical to the expressions in the semiclassical literature [1, 7, 5], once $G_{j\alpha\gamma}$ evaluated in its equivalent sum-over-states form:
$G_{j\alpha\gamma}=-2\hbar\,\mbox{Re}\sum_{j^{\prime}\neq j}\frac{\langle u_{j{\bf k}}|\,v_{\alpha}\,|u_{j^{\prime}{\bf k}}\rangle\langle u_{j^{\prime}{\bf k}}|\,v_{\gamma}\,|u_{j{\bf k}}\rangle}{(\epsilon_{j{\bf k}}-\epsilon_{j^{\prime}{\bf k}})^{3}}.$ (23)
[…]
When all three terms are included, the quadratic Hall conductivity in the $\omega$-domain is
$\sigma_{\alpha\beta\gamma}(\omega)$ $=$ $\frac{e^{3}}{\hbar}\sum_{j}\int_{\rm BZ}\frac{d{\bf k}}{(2\pi)^{d}}\;\theta(\epsilon_{\rm F}-\epsilon_{j{\bf k}})$
$\times$ $\bigg[\partial_{k_{\alpha}}G_{j\beta\gamma}-\partial_{k_{\beta}}G_{j\alpha\gamma}$
$+$ $\frac{1}{\hbar}\partial_{k_{\gamma}}\tilde{\Omega}_{j}(k_{\alpha},k_{\beta})\frac{i}{\omega+i\eta}$
$+$ $\frac{1}{\hbar^{2}}\frac{\partial^{3}\epsilon_{j{\bf k}}}{\partial k_{\alpha}\partial k_{\beta}\partial k_{\alpha}}\left(\frac{i}{\omega+i\eta}\right)^{2}\bigg],$
where $\tilde{\Omega}_{j}(k_{\alpha},k_{\beta})$ is the Berry curvature of band $j$ [19].
Finally, when the infinitesimal $\eta$ is heuristically replaced with the inverse of a relaxation time $\tau$ one gets the semiclassical result, in the form reported e.g. in Ref. [5]. The zero-temperature Fermi function $\theta(\epsilon_{\rm F}-\epsilon)$ can also be heuristically replaced by its finite-temperature counterpart.
Discussion.—I have shown that both terms of the intrinsic Hall conductivity have a geometrical formulation beyond the Bloch setting, which also applies to a larger class of materials: those where disorder and interaction play an important role. When the system is crystalline the theory is formulated à la Kohn-Sham; the resulting expressions coincide with the semiclassical ones in the $\tau\rightarrow\infty$ limit and for zero temperature. This agrees with the common wisdom that the semiclassical approximation is exact when addressing dc transport properties in a crystalline system of non interacting electrons. In this context, the semiclassical approximation is not an approximation, after all.
The major result of this work is the exact many-body expression for $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$, given as the curl in the $\kappa$ variable (the flux) of a positional-shift tensor: a Berry curvature whose variables are the flux and the electric field, having an equivalent sum-over-states expression. Hybrid geometrical quantities of the same family are at the heart of polarization theory [19]. This tensor yields indeed the linear static polarizability of a metal, where the divergent Drude term has been discounted.
When the system is cristalline the main entry of $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$ is an analogous band curvature whose variables are instead the Bloch vector and the electric field. The equivalent sum-over-states form coincides with what one finds in the literature [1, 7, 5], while the compact curvature form is possibly new. The latter, as explained above, could be computationally more appealing.
Finally, a short digression about the extrinsic effects. In the case of the linear Hall conductivity they come in two kinds: $\tau^{0}$ (called “side-jump”) and $\tau^{1}$ (called “skew scattering”) [28], while the terms “intrinsic” and “geometrical” are used as synonymous (the Bloch geometry of the pristine crystal is intended). When the disordered system is addressed by means of a supercell (ideally in the $L\rightarrow\infty$ limit) the effects of disorder become by construction intrinsic; we previously argued in Ref. [29] that the geometrical conductivity of the disordered system includes the side-jump contributions, besides the pristine-crystal geometrical response. A similar statement holds for the quadratic Hall conductivity: the $\tau^{0}$ extrinsic effects become intrinsic and included in the present generalized geometrical formulation of $\sigma^{(\rm ps)}_{\alpha\beta\gamma}$.
Acknowledgments.—I am deeply indebted to Ivo Souza for the many invaluable conversations we had about this topic. Work supported by the Office of Naval Research (USA) Grant No. N00014-20-1-2847.